# On the relation between fractional charge and statistics

### v2

@@ Abstract
We revisit an argument, originally given by Kivelson and Roček, for why the existence of fractional charge necessarily implies fractional statistics. In doing so, we resolve a contradiction in the original argument, and in the case of a $\nu=1/m$ Laughlin holes, we also show that the standard relation between fractional charge and statistics is necessary by an argument based on a t’Hooft anomaly in a one-form global ${\mathcal{Z}}_{m}$ symmetry.
In the early days of quantum Hall physics, the connection between quantized Hall conductance, fractional charge, and fractional statistics was by no means obvious. One of the landmarks in the understanding of the fractional quantum Hall effect was Laughlin’s 1983 argument [1] for why a quantized quantum Hall conductance implies fractional charge. A likewise early, but less known, argument for why fractional charge implies fractional statistics was given in 1985 by Kivelson and Roček [2]. Importantly, their argument applies more generally than to just quantum Hall systems, since it relies on only minimal theoretical assumptions. These arguments marked the beginning of the exploration of topological quantum matter.
The contemporary understanding of the topological properties of quantum Hall fluids is in terms of topological quantum field theories, typically involving Chern-Simons terms [3] and closely connected conformal field theories [4]. There is also a well-established connection to the mathematics of braided tensor categories; for a pedagogical introduction, see Ref. [5]. The extensive work on various concrete wave functions [1, 6, 7] also fits nicely into these general theoretical frameworks, while simultaneously providing essential information about the excitations and edge dynamics of many observed quantum Hall states.
@@
The original argument by Kivelson and Roček was later developed by Karlhede, Kivelson and Sondhi[8], and we now recapitulate their version. Assume that we have two quasiparticles with fractional charge $\nu e$, which we shall refer to as anyons, and imagine threading a thin unit flux tube through one of them. Next, move the other particle adiabatically around the flux-charge composite. This process will result in an Aharonov-Bohm (AB) phase $2\pi\nu$. But since the system is made entirely out of charge $e$ electrons, a unit flux is invisible; the Byers-Yang theorem [9]. Thus, there must be a compensating $-2\pi\nu$ (mod $2\pi$) braiding phase, and thus a $-\nu\pi$ exchange phase. This is the statistical phase. Note that this argument does not evoke the physics of a quantum Hall liquid. Upon closer examination, one realizes that this argument cannot be correct. Suppose that we insert $n$ fluxes instead of just one, then we would conclude that the statistical phase is $-n\nu\pi$ not $-\nu\pi$.
@@
[…]
Now consider a correlation function $\langle WB_{\nu}(C_{1})W_{\nu}(C_{2})\rangle=e^{2\pi\nu L(C_{1},C_{2})i}$ , where $L(C_{1},C_{2})$ is the linking number of the two loops[11]. This phase is just the AB phase picked up when the $\nu e$ charge encircles the unit flux vortex. It is now clear that we cannot just remove the flux since that would change the value of a gauge invariant correlation function. On the other hand, removing a unit flux in a system consisting entirely of charge $e$ electrons is a regular (global) gauge transformation. Thus, to remove the second flux to get a description entirely in terms of fractionally charged quasiparticles, we must add a compensating statistical phase to restore the original value of the correlation function. Note that the sign of this phase is opposite to the one in the original Kivelson-Roček argument. This obstruction to a regular global gauge transformation is reminiscent of a global anomaly. In this case, however, it is not the partition function that acquires an anomalous phase, but rather certain gauge invariant correlation functions related to braided loops. The connection to anomalies is, however, somewhat subtle and will be discussed below. An argument very similar in spirit to the above was given earlier by Feldman and Halperin[12].
@@
To give a formal version of the above heuristic argument, we consider two kinds of charges, which we refer to as charge and vortex, respectively, with the corresponding currents $j$ and $j_{v}$, where the former couples to the electromagnetic field $A$. The AB phase related to moving a charge $q$ around an $n$-vortex is encoded in the following BF Lagrangian,
${\cal L}_{AB}=\frac{1}{\pi}adb-qaj-nbj_{v}-eqAj,$ (1)
where we also added the coupling to the external field. Next, consider a composite of a charge $q$ and $n$ vortices, i.e. put $j_{v}=j$; we shall comment on this identification later. Introducing the new variables
$\tilde{a}$ $=nb+qa$ (2)
$\tilde{b}$ $=nb-qa\,,$
we can rewrite (1) as
${\cal L}_{AB}=\frac{1}{4\pi qn}\tilde{a}d\tilde{a}-\tilde{a}j-eqAj$ (3)
where we neglected a boundary term and also the term $\sim\tilde{b}d\tilde{b}$ since $\tilde{b}$ does not couple to the current. Specializing to a Laughlin hole obtained by inserting one flux quantum we have $q=1/m$ and $n=1$, gives
${\cal L}_{1/m}=\frac{m}{4\pi}\tilde{a}d\tilde{a}-\tilde{a}j-\frac{e}{m}Aj,$ (4)
which is the hydrodynamic action for a $\nu=1/m$ Laughlin state, known to give a statistical interaction to the sources with a statistical exchange parameter $\theta=\nu\pi$[13]. Thus, we have shown that the fundamental, i.e. the charge $e/m$, Laughlin holes created by insertion of a unit flux can alternatively be described as fractionally charged point-like anyons where the fluxes are removed.
At first sight, this result looks strange. We started from two composite particles, each with a unit flux and a charge $q=\nu e$, and thus we would expect a braiding phase $\gamma=2\times 2\pi\nu=4\pi\nu$, while (4) gives $\gamma=2\pi\nu$, which is the correct result; fermions have $\nu=1$, and gives a unit braiding phase i.e. a sign under exchange. This apparent contradiction is resolved by noticing that the identification $j_{v}=j$ changes the configuration space, and such identifications will indeed change statistics. The most famous example is given in the original article on fractional statistics by Leinaas and Myrheim[14], and for the present case, the explanation is given in Ref. [15]. Related arguments for the absence of the factor of 2 in a somewhat different context is given in Ref. [16].
Note that in the heuristic argument, which does not involve any identification of currents, we cannot start from an assembly of charges threaded by fluxes, but must proceed stepwise: Starting from the already explained two-body case, we can move to three particles by introducing a new flux-charge composite in the presence of the two $e/m$ charges obeying $\pi/m$ fractional statistics. We can then again remove the flux at the expense of introducing fractional statistics between the new charge and the two already present. Clearly, this way of reasoning can be extended to any number of particles, but it is not clear how such a recursive argument could be formulated rigorously. To do so, we return to the formal argument based on Eqs. (14).
For $n\neq 1$, we first note that taking $n=m$ gives a trivial braiding phase as expected, while $1<n<m$ amounts to having a non-integer level number, signaling that the theory is not consistent. In fact, it gives a braiding phase $\gamma=2\pi\nu n$, which depends on $n$ contrary to the Byers-Yang theorem. However, if we insert $n$ fluxes in a $\nu=1/m$ Laughlin liquid we get the charge $q=n/m$, so after the rescaling $\tilde{a}\rightarrow n\tilde{\alpha}$ (3) becomes,
${\cal L}_{1/m}=\frac{m}{4\pi}\tilde{a}d\tilde{a}-n\tilde{a}j-\frac{ne}{m}Aj$ (5)
which is consistent, and give a braiding phase $\gamma=2\pi\nu n^{2}$ as expected for a composite of $n$ anyons with $q^{\star}=e/m$. Note that in this case, different $n$ describes different particles, and it should be clear that these results again are in full agreement with our initial heuristic arguments.
@@
It remains to understand why it is not consistent to remove the flux from a composite of one charge and several fluxes, and we now show that such a theory is anomalous. For this, we first notice that taking $n>1$ in (3) we have a $U(1)$ CS theory with a fractional level number, and when defined on a closed manifold, such a theory has a global $U(1)$ anomaly. (For a simple explanation, see Ref. [17].) This does, however, not fully reflect the heuristic arguments given at the beginning of this article, where there was no need to assume a closed manifold. We now give an alternative way to understand the anomaly that works also for an open manifold. To do so, return to (3) and set the charge $q=1/m$,
${\cal L}_{1/m}=\frac{m}{4\pi n}\tilde{a}d\tilde{a}-\tilde{a}j-\frac{e}{m}Aj\,.$ (6)
@@
Following Ref. [18], we note that this theory has a global one-form ${\mathcal{Z}}_{m}$ symmetry that is generated by the Wilson lines,
$U_{e^{2\pi ik/m}}(M^{(1)})=\exp\left(ik\oint_{M^{(1)}}\tilde{a}\right)$ (7)
@@
where $M^{(1)}$ is an one-manifold, and $k=1\dots m-1$. The Wilson loops corresponding to the source $j$ in (6) are themselves generators of this algebra, and as stressed in Ref. [18], they are also charged under the symmetry $U_{e^{2\pi k/m}}(M^{(1)})$ since from (6) it follows that
$U_{e^{2\pi i{k}/m}}(M^{(1)})\exp\left(i\oint_{{\mathcal{C}}^{(1)}}\tilde{a}\right)=e^{\frac{2\pi ikn}{m}}\exp\left(i\oint_{{\mathcal{C}}^{(1)}}\tilde{a}\right)\,,$ (8)
@@
where $M^{(1)}$ is a small loop around ${\mathcal{C}}^{(1)}$. For $n=1$, $U_{e^{2\pi ik/m}}(M^{(1)})$ generates the ${\mathcal{Z}}_{m}$ symmetry corresponding to the $m$ ground states on a torus. In the case of $n$ fluxes with $n>1$, we now ask whether or not they can be removed with impunity. If they can, we should get the same result for any correlation function of the charged Wilson loops if we insert any number of $U_{e^{2\pi k/m}}(M^{(1)})$ and then average over $n$ for fixed $k$, which amounts to gauging the symmetry. However, from (8), we see that such an average will give zero, which is a consequence of the symmetry having a t’ Hooft anomaly as explained in Ref. [18]. From this, we conclude that for flux charge composites with more than one flux per charge, it is not possible to remove the flux, which was what we set out to prove.
@@
Concerning the internal structure of the quasiparticles: If we take the composite picture at face value and assume that the flux is separated from the charge, the situation is well understood[19, 20]; rotating the system an angle $2\pi$ amounts to braiding the flux and the charge within the composite. The corresponding phase angle $\pi/m=\pi\nu$, is precisely what is expected for particle with spin $s=\nu/2$, which is the generalized spin-statistics relation[21]. Also, in the context of CS theory, the fractional spin can be calculated as the orbital angular momentum as first shown in Ref. [22]. It is not clear, however, how these arguments can be formalized to extend the description based on (1), although a natural starting point would be to formulate that theory on a spatially curved surface.
Finally, we stress that our main result is that fractional charge necessarily implies fractional statistics as originally claimed in Refs. [2, 8]. This result is independent of assumptions about the internal state of flux-charge composites, which, however, is important when discussing the spin. In particular, we showed that starting from the BF lagrangian (1), which was written just to describe the AB phases related to charges encircling fluxes, we can, by identifying the charge and vortex currents, derive the hydrodynamic theory (3), and (5) which does not involve any fluxes, and correctly incorporates the fractional statistics between point-like anyons.

### v4

@@ Abstract
We revisit an argument, originally given by Kivelson and Roček, for why the existence of fractional charge necessarily implies fractional statistics. In doing so, we resolve a contradiction in the original argument, and in the case of a $\nu=1/m$ Laughlin holes, we also show that the standard relation between fractional charge and statistics is necessary by an argument based on a t’Hooft anomaly in a global one-form ${\mathcal{Z}}_{m}$ symmetry.
In the early days of quantum Hall physics, the connection between quantized Hall conductance, fractional charge, and fractional statistics was by no means obvious. One of the landmarks in the understanding of the fractional quantum Hall effect was Laughlin’s 1983 argument [laughlin1983anomalous] for why a quantized quantum Hall conductance implies fractional charge. A likewise early, but less known, argument for why fractional charge implies fractional statistics was given in 1985 by Kivelson and Roček [kivelson1985]. Importantly, their argument applies more generally than to just quantum Hall systems, since it relies on only minimal theoretical assumptions. These arguments marked the beginning of the exploration of topological quantum matter.
The contemporary understanding of the topological properties of quantum Hall fluids is in terms of topological quantum field theories, typically involving Chern-Simons terms [wen1995topological] and closely connected conformal field theories [MOORE1991362]. There is also a well-established connection to the mathematics of braided tensor categories; for a pedagogical introduction, see Ref. [bais2009condensate]. The extensive work on various concrete wave functions [laughlin1983anomalous, jain2007composite, hansson2017quantum] also fits nicely into these general theoretical frameworks, while simultaneously providing essential information about the excitations and edge dynamics of many observed quantum Hall states.
@@
The original argument by Kivelson and Roček was later developed by Karlhede, Kivelson and Sondhi[karlhede1993], and we now recapitulate their version. Assume that we have two quasiparticles with fractional charge $\nu e$, which we shall refer to as anyons, and imagine threading a thin unit flux tube through one of them. Next, move the other particle adiabatically around the flux-charge composite. This process will result in an Aharonov-Bohm (AB) phase $2\pi\nu$. But since the system is made entirely out of charge $e$ electrons, a unit flux is invisible; the Byers-Yang theorem [byersyang61]. Thus, there must be a compensating $2\pi\nu$ (mod $2\pi$) braiding phase, and thus a $\nu\pi$ exchange phase. This is the statistical phase. Note that this argument does not evoke the physics of a quantum Hall liquid. Upon closer examination, one realizes that this argument cannot be correct. Suppose that we insert $n$ fluxes instead of just one, then we would conclude that the statistical phase is $n\nu\pi$ not $\nu\pi$.
@@
First, we recall a version of Laughlin’s argument for fractional charge. If we adiabatically thread a thin flux through a $\sigma_{H}=\nu\,\sigma^{0}_{H}$ FQH state, the induced electromotive force will push the liquid radially outwards. For $n$ unit fluxes, $n\phi_{0}$, the Hamiltonian is back to its original form up to a regular gauge transformation. This means that the flux can be removed, and we are left with a localized fractional charge; the compensating fraction being pushed to the edge of the sample. Next, repeat the procedure to insert another charge flux composite. We now show that there is an obstruction against removing this second flux. For this, note that the world line of the charge-flux composite can be thought of as a t’ Hooft and a Wilson loop superimposed on each other, $WB_{\nu}(C)\equiv W_{\nu}(C)B(C)$; such superpositions have been considered earlier by Itzhaki[itzhaki2003]. Here, $W(C)=\exp[i\oint_{C}dX^{i}A_{i}(x)]$, and the t’Hooft loop, $B(C)$, implements a gauge transformation that has a $2\pi$ phase jump at an area spanned by the curve $C$.
Now consider a correlation function $\langle WB_{\nu}(C_{1})W_{\nu}(C_{2})\rangle=e^{2\pi\nu L(C_{1},C_{2})i}$ , where $L(C_{1},C_{2})$ is the linking number of the two loops[thooft81]. This phase is just the AB phase picked up when the $\nu e$ charge encircles the unit flux vortex. It is now clear that we cannot just remove the flux since that would change the value of a gauge invariant correlation function. On the other hand, removing a unit flux in a system consisting entirely of charge $e$ electrons is a regular (global) gauge transformation. Thus, to remove the second flux to get a description entirely in terms of fractionally charged quasiparticles, we must add a compensating statistical phase to restore the original value of the correlation function. Note that the sign of this phase is opposite to the one in the original Kivelson-Roček argument. This obstruction to a regular global gauge transformation is reminiscent of a global anomaly. In this case, however, it is not the partition function that acquires an anomalous phase, but rather certain gauge invariant correlation functions related to braided loops. The connection to anomalies is, however, somewhat subtle and will be discussed below. An argument very similar in spirit to the above was given earlier by Feldman and Halperin[feldman2021].
@@
Note that the above argument still does not provide a unique answer for the statistical angle, since the argument could equally well be made for the removal of more than one flux quantum. To understand the inconsistency with the removal of several flux quanta, we formulate the Kivelson-Roček argument in the language of Chern-Simons Lagrangians. The statistical interaction of a collection of charged particles and vortices is modeled by the BF Lagrangian
${\cal L}_{AB}=\frac{1}{2\pi}adb-qaj-nbj_{v}-qeAj,$ (1)
which encodes a coupling to an external electromagnetic field $A$, and the Aharonov-Bohm phases obtained by braiding a charge $q$ particle around vortices with $n$ flux quanta.
The Kivelson-Roček argument can now be stated as follows. Imagine that we consider a current of a charge-vortex composite $j_{c}$, which will have an Aharonov-Bohm phase $\theta=2\pi qn$ both when encircling a particle and a vortex. For $q=p/m$, with $p$ and $m$ relatively prime, and $n$ not a multiple of $m$, the braiding phase is fractional, meaning that an integer-valued flux is observable, in contradiction to the Byers-Yang theorem. We are thus forced to add a compensating statistical interaction between the particles, by a Chen-Simons term, so the resulting model for the particles only becomes,
${\cal L}_{CS}=-\frac{1}{4\pi qn}ada-aj-qeAj\,.$ (2)
Note that taking the values for the original Kivelson-Roceck argument, ie. $n=1$ and $q=1/m$ we retain the hydrodynamic action for a $\nu=1/m$ Laughlin state, known to give a statistical interaction to the sources with a statistical exchange parameter $\theta=\nu\pi$[wen1992classification].
@@
We now demonstrate why it is inconsistent to remove the flux from a composite of one charge and several fluxes. First notice that again taking $q=1/m$, but also $n>1$ in (2) is a $U(1)$ CS theory with a fractional level number. When defined on a closed manifold, such a theory has a global $U(1)$ anomaly. (For an explanation, see e.g. Ref. [arouca2022quantum].) This does, however, not fully reflect the heuristic arguments given at the beginning of this article, where there was no need to assume a closed manifold. We now give an alternative way to understand the anomaly that works also for an open manifold. To do so, we again put $q=1/m$ in (2),
${\cal L}_{1/m}=\frac{m}{4\pi n}ada-aj-\frac{e}{m}Aj\,.$ (3)
@@
Following Ref. [gaiotto2015generalized], we note that this theory has a global one-form ${\mathcal{Z}}_{m}$ symmetry that is generated by the Wilson lines,
$U_{e^{2\pi ik/m}}(M^{(1)})=\exp\left(ik\oint_{M^{(1)}}a\right)$ (4)
@@
where $M^{(1)}$ is an one-manifold, and $k=1\dots m-1$. The Wilson loops corresponding to the source $j$ in (3) are themselves generators of this algebra, and as stressed in Ref. [gaiotto2015generalized], they are also charged under the symmetry $U_{e^{2\pi k/m}}(M^{(1)})$ since from (3) it follows that
$U_{e^{2\pi i{k}/m}}(M^{(1)})\exp\left(i\oint_{{\mathcal{C}}^{(1)}}a\right)=e^{\frac{2\pi ikn}{m}}\exp\left(i\oint_{{\mathcal{C}}^{(1)}}a\right)\,$, (5)
@@
where $M^{(1)}$ is a small loop around ${\mathcal{C}}^{(1)}$. For $n=1$, $U_{e^{2\pi ik/m}}(M^{(1)})$ generates the ${\mathcal{Z}}_{m}$ symmetry corresponding to the $m$ ground states on a torus. In the case of $n$ fluxes with $n>1$, we now ask whether or not they can be removed with impunity. If they can, we should get the same result for any correlation function of the charged Wilson loops if we insert any number of $U_{e^{2\pi k/m}}(M^{(1)})$ and then average over $n$ for fixed $k$, which amounts to gauging the symmetry. However, from (5), we see that such an average will give zero, which is a consequence of the symmetry having a t’ Hooft anomaly as explained in Ref. [gaiotto2015generalized]. From this, we conclude that for flux charge composites with more than one flux per charge, it is not possible to remove the flux, which was what we set out to prove.
@@
Can the above argments can be generalized to filling fractions $\nu=p/q$, with $q=2n+1$? In this case the level factor of the relevant CS term will be $q/np$, so even for $n=1$, the level factor is fractional since $p$ and $q$ are relative prime. (Note that this is true even for $q=1$, corresponding to an integer number of filled Landau levels.) This leads to the conclusion that we cannot use a Kivelson-Roček argument to obtain a description of the filling fractions $\nu=p/q$ using only a single CS field. This is expected, since all Wen-Zee hydrodynamic theories for these filling fractions are based on a generalized multi-level Abelian CS theory,
${\cal L}_{WZ}=\frac{1}{4\pi}K_{ij}a^{i}db^{j},$ (6)
where, for example, when $\nu=n$, $K_{ij}$ is an $n$-dimensional unit matrix.
We have so far not discussed the internal structure of the quasiparticles. To do so, we first present a simple BF theory for anyons that, in terms of braiding phases and coupling to the external electromagnetic field, is equivalent to the CS theory (2). The idea is to model the anyons as charge-flux composites, and we now show that this indeed leads to a theory equivalent to the CS theory. The starting point is
${\cal L}_{AB}=\frac{1}{\pi}adb-qaj-nbj_{v}-qeAj,$ (7)
where now $j_{v}$ is the current of half-vortices (as the physical vortices in a superconductor). Next, introduce the two fields $a_{\pm}=qa\pm nb$ and use the identity $4nqbda=a_{+}da_{+}-a_{-}da_{-}$, to rewrite (7) as
${\cal L}_{AB}=\frac{1}{4\pi qn}(a_{+}da_{+}-a_{-}da_{-})-\frac{1}{2}a_{+}(j+j_{v})-\frac{1}{2}a_{-}(j-j_{v})-qeAj\,{.}$ (8)
Now we take the flux and charge to be combined in a composite, described by the current $j_{c}$, meaning that the flux and charge currents are not independent, $j_{c}=j=j_{v}$. Consequently, the field $a_{-}$ decouples, and we are left with
${\cal L}_{c}=\frac{1}{4\pi qn}a_{+}da_{+}-a_{+}j_{c}-qeAj_{c}\,$ (9)
which is the Lagrangian (2).  [^1] Note that these composites are not the same as the “cyons” discussed in Ref. [cyons], which are charge-flux composites interacting by a Maxwell term. If we take this model of the anyons as charge-flux composites at face value and assume that the flux is separated from the charge, there is a well understood physical consequence [PhysRevLett.48.1144, hansson1991spin]; rotating the system an angle $2\pi$ amounts to braiding the flux and the charge within the composite. The corresponding phase angle $\pi/m=\pi\nu$, is precisely what is expected for particle with spin $s=\nu/2$, which is the generalized spin-statistics relation[balachandran1990topological]. Also, in the context of CS theory, the fractional spin can be calculated as the orbital angular momentum as first shown in Ref. [hagen1985rotational].
Finally, we stress that our main result is that fractional charge necessarily implies fractional statistics as originally claimed in Refs. [kivelson1985, karlhede1993]. This result is independent of assumptions about the internal state of flux-charge composites, which, however, is important when discussing the spin. In particular, we showed that starting from the BF Lagrangian (1), which was written just to describe the AB phases related to charges encircling fluxes, we can,derive the hydrodynamic CS theory (2). We also presented a an equivalent BF description where the anyons are flux-charge composites, which carey a spin consistent with the generalized spin-statistics theorem.